Rob is planning a three hour trip down the Allenem River and back to his starting point. He knows that he can paddle in still water at 3 miles per hour and that the rate of the current is 2 miles per hour. How much time can he spend going downstream? How far dwonstream can he travel?
Answer by mananth(16946)

let x be the distance he can travel to complete trip in 3 hours
downstream speed = 3+2 = 5mph
upstream speed = 3-2 = 1mph
time downstream + time upstream = 3 hours
x/5 + x = 3
6x/5 =3
6x=15
x= 2.5 miles.
..
speed downstream = 5 mph
distance = 2.5 miles
time downstream = 5/2.5 = 2 hours
...
m.ananth@hotmail.ca